# Diversifying Query: Region-Guided Transformer for Temporal Sentence Grounding

### v1

@@ Abstract
Temporal sentence grounding is a challenging task that aims to localize the moment spans relevant to a language description. Although recent DETR-based models have achieved notable progress by leveraging multiple learnable moment queries, they suffer from overlapped and redundant proposals, leading to inaccurate predictions. We attribute this limitation to the lack of task-related guidance for the learnable queries to serve a specific mode. Furthermore, the complex solution space generated by variable and open-vocabulary language descriptions exacerbates the optimization difficulty, making it harder for learnable queries to distinguish each other adaptively. To tackle this limitation, we present a Region-Guided TRansformer (RGTR) for temporal sentence grounding, which diversifies moment queries to eliminate overlapped and redundant predictions. Instead of using learnable queries, RGTR adopts a set of anchor pairs as moment queries to introduce explicit regional guidance. Each anchor pair takes charge of moment prediction for a specific temporal region, which reduces the optimization difficulty and ensures the diversity of the final predictions. In addition, we design an IoU-aware scoring head to improve proposal quality. Extensive experiments demonstrate the effectiveness of RGTR, outperforming state-of-the-art methods on QVHighlights, Charades-STA and TACoS datasets.
@@ 1 Introduction
Temporal sentence grounding (TSG) aims at localizing the moment spans semantically aligned with the given language description in an untrimmed video. This area of research has drawn increasing attention in recent years due to its wide range of potential applications, such as human-computer interaction [30, 10] and information retrieval [22, 37]. Early methods address the TSG task by designing predefined dense proposals [6, 41, 35] or directly learning sentence-frame interactions [15, 16, 39]. The recent success of detection transformer (DETR) [2] has inspired the integration of transformers into the temporal sentence grounding framework [11, 24, 36]. By decoding moment spans from a set of learnable queries, they streamline the complicated grounding pipeline.
[…]
-0.21cm0cm Refer to caption
@@
Figure 1: Visualization of all moment predictions on QVHighlights val split, for the 3 selected moment queries in EaTR [9], TR-DETR [31], CG-DETR [23] and RGTR (Ours). x-axis denotes the normalized moment span center coordinates, y-axis denotes the normalized moment span length. It can be observed that all queries in previous methods generate numerous overlapped and redundant proposals. For example, the second query tends to predict long moments near the middle of the videos, but the proposals of short moments conflict with this purpose, leading to inefficient predictions. In contrast, the predicted region of each query in our RGTR is distinct and more concentrated.
@@
Although DETR-based approaches have achieved notable performance in TSG task, we still observe some unique limitations of the DETR structure compared to other fields (i.e., object detection and temporal action localization). Specifically, they suffer from limited query distribution and overlapped proposals, leading to inaccurate predictions. As shown in Fig. 1, we present the distribution of three representative moment queries in different methods, each query predicting different temporal regions ( e.g., the lower left area represents a short moment near the video’s start and the higher middle area represents a long moment). In previous methods EaTR [9], TR-DETR [31] and CG-DETR [23], each query includes numerous overlapped and redundant proposals for the same region (i.e., the short moments in the lower part), resulting in inefficient predictions. We attribute this limitation to the lack of task-related guidance (e.g., category constraints, spatial distribution prior, etc.) for the learnable queries to serve a specific mode. The task-related guidance has been scarcely explored in TSG task although it is crucial to improve query diversity. Furthermore, the complex solution space generated by variable and open-vocabulary language descriptions exacerbates the optimization difficulty, making it harder for learnable queries to distinguish each other adaptively. Another limitation is that the proposal scoring in previous methods is purely based on the classification confidence (foreground or background), ignoring the quality of the predicted boundary. Instead, we argue that correctly classified proposals that better overlap with the ground-truth should be assigned higher scores. The above limitations significantly restrict the accurate localization of the DETR structure in TSG task.
@@
In this paper, we introduce an effective Region-Guided TRansformer (RGTR) framework to cope with the aforementioned limitations in the TSG task. To address the issue of limited query diversity, we introduce regional priors based on the distribution of ground-truth moment spans as task-related guidance. Specifically, we design a region-guided decoder with a new concept of anchor pairs as moment queries to provide regional guidance. Each anchor pair consists of a static anchor and a dynamic anchor, both of which are initialized by different clustering points on the ground-truth moment spans. Such explicit initialization imposes the regional priors as guidance on each anchor pair, enhancing the diversity of the query distributions. The two types of anchors serve different roles in the decoder, collaboratively guiding localization with explicit regional guidance and eliminating overlapped predictions. In addition, to improve the scoring of high-quality proposals, we propose an IoU-aware scoring head. By supervising the IoU score with L2 loss, the prediction head considers both classification confidence and localization quality. As shown in Fig. 1, our RGTR eliminates redundant predictions and exhibits diverse query distributions compared to previous methods.
@@
Our main contributions are summarized as follows. (1) We design a novel region-guided decoder, which adopts a set of explicitly initialized anchor pairs as moment queries to introduce regional priors as task-related guidance. (2) We propose an IoU-aware scoring head that incorporates localization quality to enhance classification confidence estimation and distinguish high-quality proposals. (3) By employing these techniques, we introduce a Region-Guided TRansformer that diversifies the moment queries and improves proposal quality. Extensive experiments conducted on three challenging benchmarks demonstrate the superiority of the proposed model over state-of-the-art methods.
2 Related Works
@@
Temporal Sentence Grounding in Videos.  Temporal sentence grounding aims at predicting the moment spans of the described activity given an untrimmed video and a language description, which is first proposed in [6, 1]. Early methods fall into two main categories: proposal-based methods and proposal-free methods. Proposal-based methods [8, 17, 42, 35] initially generate multiple candidate proposals and rank them based on their similarity with the description. Despite achieving promising performance, these methods are greatly limited by the high computational cost of proposal matching. Proposal-free methods [3, 21, 16, 39] are proposed to avoid the need for predefined candidate moments. Instead of relying on segment candidates, they directly predict the start and end boundaries of the target moments by leveraging cross-modal interactions between video and sentence.
The recent success of detection transformer (DETR) [2] has inspired the integration of transformers into the temporal sentence grounding framework [11, 31, 24]. DETR-based methods simplify the whole process into an end-to-end manner by removing handcrafted techniques. However, due to the lack of task-related guidance for the learnable queries to serve a specific mode, almost all the previous methods generate numerous overlapped and redundant predictions. In contrast, our RGTR improves the diversity of queries and eliminates redundant predictions by introducing explicit regional priors.
Detection Transformers.  Recently, Transformer [32] has raised great attention in computer vision [4, 33, 2]. The adoption of transformers to object detection (DETR) [2] builds a fully end-to-end object detection system based on transformers, which largely simplifies the traditional detection pipeline. It also achieves notable performances compared with highly optimized CNN-based detectors.
@@
The formulation of decoder queries has been widely studied in previous works [44, 18, 28]. Anchor DETR [34] initializes queries based on anchor points for specific detection modes. DAB-DETR [18] formulates decoder queries with content and action embeddings. DINO [40] adds position priors for the positional query and randomly initializes the content query. Motivated by their great success, we introduce a set of anchor pairs to introduce the explicit regional priors for accurate prediction.
@@ 3 Method
In this section, we briefly present an overview of our proposed framework in Sec. 3.1. Then we elaborate on the main components of RGTR, including cross-modal alignment encoder (in Sec. 3.2), region-guided decoder (in Sec. 3.3) and IoU-aware scoring head (in Sec. 3.4). Finally, the training objectives are introduced in Sec. 3.5.
{adjustwidth}
-0.2cm0cm Refer to caption
Figure 2: Overview of the proposed RGTR architecture. Given a video and a text description, we first employ two frozen pre-trained models to extract visual and textual features. Subsequently, the cross-modal alignment encoder is constructed to align and fuse the visual and textual features effectively. Then, we design a region-guided decoder to introduce the regional priors through a set of explicitly initialized anchor pairs. Finally, the IoU-aware scoring head generates high-quality moment spans by incorporating localization quality to enhance the classification confidence estimation.
@@ 3.1 Overview
Given an untrimmed video $\mathcal{V}=\left\{{v_{t}}\right\}^{L}_{t=1}$ with $L$ frames and an associated natural language description $\mathcal{T}=\left\{{t_{n}}\right\}^{N}_{n=1}$ with $N$ words, Temporal Sentence Grounding (TSG) aims to accurately predict the moment span $\hat{m}=(\hat{m}_{c},\hat{m}_{\sigma})$ that is most relevant to the given text description, where $\hat{m}_{c}$ and $\hat{m}_{\sigma}$ represent the center time and duration length of the moment span.
Recent DETR-based methods [11, 24, 38] replace hand-crafted components with learnable positional queries to predict target moments. These positional queries, representing a set of learnable referential search areas, are initialized as random learnable embeddings in the previous methods [24, 38, 36]. However, due to the lack of task-related guidance (e.g., categories constraints, spatial distribution prior, etc.) and the extensive variability of language descriptions, the random initialization of positional queries greatly exacerbates the optimization difficulty.
To address this problem, we propose a Region-Guided TRansformer (RGTR), which adopts a set of explicitly initialized anchor pairs to replace randomly initialized learnable queries without explicit guidance. In our framework, we construct a region-guided decoder through anchor pairs to provide directive and diverse reference search areas. In addition, we introduce an IoU-aware scoring head to distinguish high-quality proposals. The overall architecture is shown in Fig. 2.
@@ 3.2 Cross-Modal Alignment Encoder
Following previous methods [24, 31, 13], we use the pre-trained VGG model [29], CLIP [26] and Slowfast model [5] to extract clip-level visual features $F_{v}\in\mathbb{R}^{L\times d_{v}}$, where $L$ represents the number of clips and $d_{v}$ is the dimension of the visual features. Furthermore, we utilize the pre-trained Glove model [25] and CLIP model to extract word-level textual features $F_{t}\in\mathbb{R}^{N\times d_{t}}$, where $N$ is the number of words and $d_{t}$ is the dimension of the textual features.
Given the clip-level visual features $F_{v}$ and the word-level textual features $F_{t}$, they are first projected into the common multimodal space using multi-layer perceptrons (MLPs) to produce the corresponding features $\bar{F}_{v}\in\mathbb{R}^{L\times D}$ and $\bar{F}_{t}\in\mathbb{R}^{N\times D}$, where $D$ is the embedding dimension. As highlighted in previous work [12], aligning modalities before interaction could reduce the modal gap and obtain better modal representations. Therefore, we employ an alignment loss $\mathcal{L}_{\text{align}}$ to facilitate the alignment between videos and sentences by using their global representations.
@@
where $B$ represents the batch size, $G^{i}_{v}\in\mathbb{R}^{D}$ and $G^{i}_{t}\in\mathbb{R}^{D}$ denote the global feature of the $i$-th video and the $i$-th sentence in a training batch, respectively. Additionally, we employ a visual feature refinement module [31] to suppress the interference of visual information irrelevant to the sentence.
After alignment, we adopt the text-to-video encoder [24] to obtain text-aware video representations. Specifically, three cross-attention layers are utilized to integrate textual features into the visual features, where query features are extracted from the visual features: $Q_{v}=\text{Linear}_{q}(\bar{F}_{v})$, key and value features are obtained from the textual features $K_{t}=\text{Linear}_{k}(\bar{F}_{t})$ and $V_{t}=\text{Linear}_{v}(\bar{F}_{t})$:
$\hat{F}_{v}=\text{Attention}(Q_{v},K_{t},V_{t})=\text{Softmax}(\frac{Q_{v}K^{\top}_{t}}{\sqrt{D}})V_{t}\in\mathbb{R}^{L\times D}.$ (2)
Subsequently, three self-attention layers are leveraged to enhance the representations to help the model better understand the video sequence relations. Here, we project $\hat{F}_{v}$ to $Q_{\hat{v}}$, $K_{\hat{v}}$ and $V_{\hat{v}}$ and use them to obtain the final cross-modal fusion embedding $F$, which is imposed by saliency score constraints $\mathcal{L}_{\text{sal}}$ [31]. Refer to the supplemental material for detailed information about $\mathcal{L}_{\text{sal}}$.
@@ 3.3 Region-Guided Decoder
Given the cross-modal fusion embedding $F$, we aim to localize moment spans semantically aligned with the language description. As discussed in Sec. 3.1, previous methods employ randomly initialized learnable queries without explicit guidance, leading to increasing optimization difficulty and reduced query diversity. In contrast, we design a region-guided decoder, which adopts a set of explicitly initialized anchor pairs as moment queries to provide directive and diverse regional priors. Each anchor pair consists of a static anchor and a dynamic anchor, both of which are initialized by different clustering points on the ground-truth moment spans. The two anchors serve different roles in the decoder, collaboratively guiding localization with explicit regional guidance. The structure of the region-guided decoder is described in Fig. 3. We elaborate on the detailed process in the following.
Figure 3: The structure of our proposed region-guided decoder with anchor pair (static anchor and dynamic anchor).
Explicit anchor initialization.  Due to the specificity of the TSG task, we lack the task-related guidance (e.g., category constraints) present in other detection tasks. Nonetheless, we can still provide regional guidance for the decoding process by considering the distribution of ground-truth moment spans. Specifically, we first modify the 2D anchor boxes in [18] to represent 1D moments in the video, i.e., the center coordinate $m_{c}$ and the duration $m_{\sigma}$ of the moments are utilized to represent the static anchor and dynamic anchor. Then, we generate $\mathcal{K}$ representative points $A\in\mathbb{R}^{\mathcal{K}\times 2}$ by adopting k-means clustering algorithm on the ground-truth of all the moment spans $m=(m_{c},m_{\sigma})$ in the training dataset. These clustering points represent explicit temporal regions with diverse center coordinates and durations. Since events described in the sentence can occur anywhere in the video, generating diverse temporal regions as guidance is crucial. Therefore, the static and dynamic anchors are initialized to $\mathcal{K}$ representative anchors: $A^{0}_{s}=A^{0}_{d}=A\in\mathbb{R}^{\mathcal{K}\times 2}$, and the positional embeddings of anchor pairs are generated by:
@@
where PE($\cdot$) means positional encoding to generate sinusoidal embeddings. For clarity, we use $A^{j}_{s}$ and $P^{j}_{s}$ to sign the static anchor in $j$-th decoder layer, even though they are never updated. With the explicit initialization, regional priors are introduced to guide different anchor pairs in generating non-overlapped predictions.
Anchor pair update.  Although introducing regional guidance by explicit initialization, maintaining the guidance during decoding iterations is also important. Following this idea, the static anchor is designed to maintain guidance without updating, while the dynamic anchor is designed to update for localization. For static anchors, we have
@@
Given the dynamic anchors $A^{j}_{d}=(a_{c}^{j},a_{\sigma}^{j})$ in $j$-th decoder layer and the relative positions $\Delta A^{j}_{d}$ = ($\Delta a^{j}_{c},\Delta a^{j}_{\sigma}$) by a prediction head, the dynamic anchors and the positional embeddings are updated as:
$A^{j+1}_{d}=A^{j}_{d}+\Delta A^{j}_{d}=(a_{c}^{j}+\Delta a^{j}_{c},a_{\sigma}^{j}+\Delta a^{j}_{\sigma}),\;\;P^{j+1}_{d}=\text{MLP}(\text{PE}(A^{j+1}_{d})).$ (5)
Note that all prediction heads in different decoder layers share the same parameters.
Region-guided attention module.  Similar to the general decoder, our region-guided decoder also includes two parts: self-attention module and cross-attention module. However, we employ different anchors in the two modules for varying roles, as shown in Fig. 3. In the self-attention module, we utilize static anchors to focus content embeddings on our preset representative temporal regions and share information across different regions, such that the updated content embedding $C_{s}^{j}$ is as follows:
$C_{s}^{j}=\text{Attention}(\mathrm{Q}:C^{j-1}+P^{0}_{s},\;\mathrm{K}:C^{j-1}+P^{0}_{s},\;\mathrm{V}:C^{j-1})\in\mathbb{R}^{\mathcal{K}\times D},$ (6)
where $C^{j-1}\in\mathbb{R}^{\mathcal{K}\times D}$ is the content embedding from ($j-1$)-th decoder layer, and $C^{0}$ is initialized to zeros. In the cross-attention module, we employ dynamic anchors to aggregate region-specific features from cross-modal fusion embedding $F$ with the assistance of region-guided content embedding $C_{s}^{j}$. Therefore, the content embedding $C^{j}$ is updated as:
$C^{j}=\text{Attention}(\mathrm{Q}:\text{Cat}(C^{j}_{s},P_{d}^{j}),\;\mathrm{K}:\text{Cat}(F,\text{PE}(F)),\;\mathrm{V}:F)\in\mathbb{R}^{\mathcal{K}\times D},$ (7)
where Cat($\cdot$) means concatenation function. By adopting anchor pairs with regional priors, the region-guided decoder ensures the diversity of the final predictions.
3.4 IoU-Aware Scoring Head
The region-guided decoder improves the quality of proposals by reducing overlapped and redundant proposals, while high-quality proposals demand not only fewer duplications but also accurate boundaries. In the previous DETR-based methods [24, 9, 31], classification confidence (foreground or background) is adopted to rank all final proposals. However, a single binary classification score may inadequately assess proposal quality by overlooking temporal boundary accuracy. To distinguish high-quality proposals, we introduce an IoU-aware scoring head, which considers both localization quality and classification confidence.
Specifically, the output of the decoder is fed to an FFN and a linear layer to predict the moment span $\hat{m}=(\hat{m}_{c},\hat{m}_{\sigma})$ and the confidence score $p_{c}$. Additionally, we add a linear layer to predict the expected IoU $p_{\text{IoU}}$ with respect to the ground-truth IoU. Instead of scoring proposals by classification confidence alone, we score them by a joint combination of confidence and IoU score, i.e., the product between $p_{c}$ and $p_{\text{IoU}}$.
We supervise the IoU score with an L2 loss to the ground-truth IoU, denoted as $\hat{g}_{\text{IoU}}$,
This additional IoU score can explicitly incorporate localization quality to enhance the classification confidence estimation, thereby generating high-quality proposals. In addition, non maximum suppression (NMS) is applied during the inference.
3.5 Training Objectives
The objective losses of RGTR include four parts: moment loss $\mathcal{L}_{\text{mom}}$, saliency loss $\mathcal{L}_{\text{sal}}$, alignment loss $\mathcal{L}_{\text{align}}$ and IoU loss $\mathcal{L}_{\text{IoU}}$. Following [24], moment loss includes L1, gIoU, and focal loss, and saliency loss includes margin ranking loss and contrastive loss. The overall objective is defined as:
where $\lambda_{\ast}$ are the balancing parameters. Refer to the supplemental material for detailed information about the training objectives.
4 Experiments
4.1 Datasets, Metrics, and Implementation Details
Datasets.  We evaluate the proposed method on three temporal grounding benchmarks, including the QVHighlights [11], Charades-STA [6], and TACoS [27] datasets. QVHighlights spans various themes from everyday lifestyle vlogs to social events in news videos. Charades-STA comprises intricate daily human activities. TACoS mainly showcases long-form videos focusing on culinary activities.
Metrics.  To make fair comparisons, we adopt the Recall@1 (R1) under the IoU thresholds of 0.3, 0.5, and 0.7 following the previous works [11, 31, 23]. Since QVHighlights contains multiple ground-truth moments per sentence, we also report the mean average precision (mAP) with IoU thresholds of 0.5, 0.75, and the average mAP over a set of IoU thresholds [0.5: 0.05: 0.95]. For Charades-STA and TACoS, we compute the mean IoU of top-1 predictions.
Table 1: Performance Comparison on QVHighlights test and val splits. We highlight the best score in each column in bold, and the second best score with underline.

### v2

@@ Abstract
Temporal sentence grounding is a challenging task that aims to localize the moment spans relevant to a language description. Although recent DETR-based models have achieved notable progress by leveraging multiple learnable moment queries, they suffer from overlapped and redundant proposals, leading to inaccurate predictions. We attribute this limitation to the lack of task-related guidance for the learnable queries to serve a specific mode. Furthermore, the complex solution space generated by variable and open-vocabulary language descriptions complicates optimization, making it harder for learnable queries to adaptively distinguish each other, leading to more severe overlapped proposals. To address this limitation, we present the Region-Guided TRansformer (RGTR) for temporal sentence grounding, which introduces regional guidance to increase query diversity and eliminate overlapped proposals. Instead of using learnable queries, RGTR adopts a set of anchor pairs as moment queries to introduce explicit regional guidance. Each moment query takes charge of moment prediction for a specific temporal region, which reduces the optimization difficulty and ensures the diversity of the proposals. In addition, we design an IoU-aware scoring head to improve proposal quality. Extensive experiments demonstrate the effectiveness of RGTR, outperforming state-of-the-art methods on three public benchmarks and exhibiting good generalization and robustness on out-of-distribution splits. Codes are available at https://github.com/TensorsSun/RGTR
@@ 1 Introduction
Temporal sentence grounding (TSG) aims at localizing the moment spans semantically aligned with the given language description in an untrimmed video. Early methods address the TSG task by designing predefined dense proposals (Gao et al. 2017; Wang et al. 2022b) or directly learning sentence-frame interactions (Liu et al. 2022a; Yang and Wu 2022). The recent success of detection transformer (DETR) has inspired the integration of transformers into the TSG framework (Moon et al. 2023b; Xiao et al. 2024). By decoding moment spans from a set of learnable queries, they streamline the complicated grounding pipeline.
@@
Figure 1: Visualization comparison of all moment predictions on QVHighlights val split, for the 3 representative moment queries in EaTR (Jang et al. 2023), TR-DETR (Sun et al. 2024), CG-DETR (Moon et al. 2023a) and RGTR (Ours). x-axis denotes the normalized moment span center coordinate, y-axis denotes the normalized moment span length. All queries in previous methods generate numerous overlapped proposals. For example, the second query tends to predict long moments near the middle of the videos (higher middle area), but the proposals of short moments (lower area) conflict with this purpose, leading to ineffective predictions. In contrast, the predicted region of each query in our RGTR is distinct and more concentrated.
@@
Although DETR-based approaches have achieved notable performance in TSG task, we still observe some unique limitations of the DETR structure compared to other fields (e.g., object detection). Specifically, they suffer from limited query distribution and overlapped proposals, leading to inaccurate predictions. As shown in Fig. 1, we present the center-length distribution of moment queries in three DETR-based methods, each query learns to predict different temporal regions (e.g., the lower left area represents a short moment near the video’s start and the higher middle area represents a long moment). In previous methods, each query includes numerous overlapped and redundant proposals for the same region (e.g., the short moments in the lower part), resulting in ineffective predictions. We attribute this limitation to the lack of task-related guidance (e.g., category constraints, spatial distribution prior, etc.) for the learnable queries to serve a specific mode. Although task-related guidance is crucial to reducing the overlapped proposals, it has been scarcely explored in TSG task. Furthermore, the complex solution space generated by variable and open-vocabulary language descriptions exacerbates the optimization difficulty, making it harder for learnable queries to distinguish each other adaptively and resulting in more severe overlapped proposals. Another limitation is that the proposal scoring in previous methods is purely based on the classification confidence, ignoring the quality of the predicted boundary. Instead, we argue that correctly classified proposals that better overlap with the ground-truth should be assigned higher scores. The above limitations significantly restrict the accurate localization of the DETR structure in TSG task.
@@
In this paper, we introduce an effective Region-Guided TRansformer (RGTR) framework to cope with the aforementioned limitations in TSG task. To address the issue of overlapped proposals, we introduce regional priors based on the distribution of ground-truth moment spans as task-related guidance. This regional guidance can eliminate overlapped proposals by increasing query diversity. Specifically, we design a region-guided decoder with a new concept of anchor pairs as moment queries to provide regional guidance. Each moment query consists of a static anchor and a dynamic anchor, both of which are initialized by different clustering centers on the ground-truth moment spans. Such explicit initialization imposes regional priors as guidance on each moment query, enhancing the diversity of query distribution. The two types of anchors serve different roles in the decoder. The static anchor is designed to maintain the regional guidance, so it is not updated during decoding. With the help of the fixed static anchor, the dynamic anchor continuously updates to make diverse predictions for various temporal regions. They collaboratively guide localization with explicit regional guidance and eliminate overlapped proposals. In addition, to improve the scoring of high-quality proposals, we propose an IoU-aware scoring head. By supervising the IoU score with L2 loss, the prediction head considers both classification confidence and localization quality.
@@
Extensive experiments on three TSG benchmarks demonstrate the effectiveness of RGTR framework. As shown in Fig. 1, RGTR eliminates redundant proposals and exhibits diverse query distributions compared to previous methods. Our main contributions are summarized as follows: (1) We design a novel region-guided decoder, which adopts a set of explicitly initialized anchor pairs as moment queries to introduce regional priors as task-related guidance. (2) We propose an IoU-aware scoring head that incorporates localization quality to enhance classification confidence estimation and distinguish high-quality proposals. (3) By employing these techniques, we introduce a Region-Guided TRansformer that eliminates overlapped proposals and improves localization quality. RGTR achieves state-of-the-art performance on three challenging benchmarks and exhibits good generalization and robustness on out-of-distribution splits.
2 Related Work
@@
Temporal Sentence Grounding.  Temporal sentence grounding aims at predicting the moment spans of the described activity given an untrimmed video and a language description, which is first proposed in (Gao et al. 2017). Early methods fall into proposal-based methods and proposal-free methods. Proposal-based methods (Liu et al. 2018; Xia et al. 2022; Wang et al. 2022b) initially generate multiple candidate proposals and rank them based on their similarity with the description. Proposal-free methods (Lu et al. 2019; Chen et al. 2020; Yang and Wu 2022) are proposed to avoid the need for predefined candidate moments. Instead of relying on segment candidates, they directly predict the start and end boundaries of the target moments. The recent success of detection transformer (DETR) (Carion et al. 2020) has inspired the integration of transformers into the temporal sentence grounding framework (Lei, Berg, and Bansal 2021; Liu et al. 2022c; Lee and Byun 2023). DETR-based methods simplify the whole process into an end-to-end manner by removing handcrafted techniques. However, due to the lack of task-related guidance for the learnable queries to serve a specific mode, almost all previous methods generate numerous overlapped and redundant proposals. In contrast, our method eliminates overlapped proposals by introducing regional guidance.
Figure 2: (a) Overview of the proposed RGTR architecture. Given a video and a text description, we first employ two frozen pre-trained models to extract visual and textual features. Subsequently, the cross-modal alignment encoder is constructed to align and fuse the visual and textual features effectively. Then, we design a region-guided decoder to introduce the regional guidance for decoding process through a set of explicitly initialized anchor pairs. Finally, the IoU-aware scoring head generates high-quality proposals by incorporating localization quality to enhance the classification confidence estimation. (b) The clustering centers with regional priors are obtained by adopting k-means algorithm on the distribution of all ground-truth moment spans.
@@
Detection Transformers.  Recently, the adoption of transformers to object detection (DETR) (Carion et al. 2020) builds a fully end-to-end object detection system based on transformers. The formulation of decoder queries has also been widely studied in previous work (Zhu et al. 2020; Shi et al. 2022, 2023). Anchor DETR (Wang et al. 2022a) initializes queries based on anchor points for specific detection modes. DAB-DETR (Liu et al. 2022b) formulates decoder queries with content and action embeddings. DINO (Zhang et al. 2022) adds position priors for the positional query and randomly initializes the content query. Motivated by their great success, we introduce a set of anchor pairs to introduce explicit regional guidance for accurate prediction.
@@ 3.1 Overview
Given an untrimmed video $\mathcal{V}=\left\{{v_{t}}\right\}^{L}_{t=1}$ with $L$ frames and an associated natural language description $\mathcal{T}=\left\{{t_{n}}\right\}^{N}_{n=1}$ with $N$ words, TSG aims to accurately predict the moment span $m=(m_{c},m_{\sigma})$ that is most relevant to the given description, where $m_{c}$ and $m_{\sigma}$ represent the center time and duration length of the moment span.
Recent DETR-based methods replace hand-crafted components with learnable positional queries to predict target moments. These positional queries, representing a set of learnable referential search areas, are initialized as random learnable embeddings in the previous methods (Moon et al. 2023b; Yang et al. 2024; Xiao et al. 2024). However, due to the lack of task-related guidance (e.g., categories constraints, spatial distribution prior, etc.) and the extensive variability of language descriptions, the random initialization of positional queries greatly exacerbates the optimization difficulty and produces numerous overlapped proposals.
To address this problem, we propose the Region-Guided TRansformer (RGTR), which adopts a set of explicitly initialized anchor pairs as moment queries to replace randomly initialized learnable queries without guidance. In our framework, we construct a region-guided decoder through anchor pairs to provide directive and diverse reference search areas for decoding process. In addition, we introduce an IoU-aware scoring head to distinguish high-quality proposals. The overall architecture is shown in Fig. 2a.
@@ 3.2 Cross-Modal Alignment Encoder
Following previous methods (Moon et al. 2023b; Li et al. 2024), we use the pre-trained CLIP (Radford et al. 2021) and Slowfast model (Feichtenhofer et al. 2019) to extract clip-level visual features $F_{v}\in\mathbb{R}^{L\times d_{v}}$, where $L$ represents the number of clips and $d_{v}$ is the dimension of visual features. Furthermore, we utilize the CLIP model to extract word-level textual features $F_{t}\in\mathbb{R}^{N\times d_{t}}$, where $N$ is the number of words and $d_{t}$ is the dimension of textual features.
Given the clip-level visual features $F_{v}$ and the word-level textual features $F_{t}$, they are first projected into the common multimodal space using multi-layer perceptrons (MLPs) to produce the corresponding features $\bar{F}_{v}\in\mathbb{R}^{L\times D}$ and $\bar{F}_{t}\in\mathbb{R}^{N\times D}$, where $D$ is the embedding dimension. As highlighted in previous work (Li et al. 2021; Sun et al. 2024), aligning modalities before interaction could reduce the modal gap and obtain better modal representations. Therefore, we employ an alignment loss $\mathcal{L}_{\text{align}}$ to facilitate the alignment between videos and sentences.
@@
where $B$ represents the batch size, $G^{i}_{v}\in\mathbb{R}^{D}$ and $G^{i}_{t}\in\mathbb{R}^{D}$ denote the global feature of the $i$-th video and the $i$-th sentence in a training batch, respectively.
After alignment, we adopt a text-to-video encoder to obtain text-aware video representations. Specifically, three cross-attention layers are utilized to integrate textual features into the visual features:
$\hat{F}_{v}=\text{Attention}(Q_{v},K_{t},V_{t})=\text{Softmax}(\frac{Q_{v}K^{\top}_{t}}{\sqrt{D}})V_{t}.$ (2)
where $Q_{v}=\text{Linear}_{q}(\bar{F}_{v})$, $K_{t}=\text{Linear}_{k}(\bar{F}_{t})$ and $V_{t}=\text{Linear}_{v}(\bar{F}_{t})$. Subsequently, three self-attention layers are leveraged to enhance the representations to help the model better understand the video sequence relations. Here, we project $\hat{F}_{v}$ to $Q_{\hat{v}}$, $K_{\hat{v}}$ and $V_{\hat{v}}$ and use them to obtain the final cross-modal fusion embedding $F$, which is imposed by saliency score constraints $\mathcal{L}_{\text{sal}}$ (Moon et al. 2023b).
@@ 3.3 Region-Guided Decoder
Given the fusion embedding $F$, we aim to localize moment spans semantically aligned with the description in the decoder. As discussed in Sec. 3.1, previous methods employ randomly initialized learnable queries without task-related guidance, leading to increasing optimization difficulty and numerous overlapped proposals. In contrast, we design a region-guided decoder, which adopts a set of explicitly initialized anchor pairs as moment queries to provide directive and diverse regional guidance. Each anchor pair consists of a static anchor and a dynamic anchor, both of which are initialized by clustering centers on the ground-truth moment spans. The two types of anchors serve different roles in the decoder, where static anchors maintain regional guidance without updating and dynamic anchors make diverse predictions. They collaboratively guide localization with explicit regional guidance. The structure of the region-guided decoder is described in Fig. 3. We elaborate on the detailed process in the following.
Anchor Explicit Initialization.  Due to the specificity of the TSG task, we lack the task-related guidance (e.g., category constraints) present in other detection tasks. Nonetheless, we can still provide regional guidance for the decoding process by considering the distribution of ground-truth moment spans. Specifically, the forms of static anchors and dynamic anchors are first defined as $a=(a_{c},a_{\sigma})$, where $a_{c}$ is the center coordinate and the $a_{\sigma}$ is the duration of the moment. Then, as shown in Fig. 2b, we generate $\mathcal{K}$ clustering centers $A\in\mathbb{R}^{\mathcal{K}\times 2}$ by adopting k-means clustering algorithm on the distribution of all ground-truth moment spans. These clustering centers represent explicit temporal regions with diverse center coordinates and durations. Since events described in the text can occur anywhere in videos, generating diverse temporal regions as guidance is crucial. Therefore, the static and dynamic anchors are initialized by $\mathcal{K}$ clustering centers: $A^{0}_{s}=A^{0}_{d}=A\in\mathbb{R}^{\mathcal{K}\times 2}$, and the positional embeddings of anchor pairs are generated by:
Figure 3: The structure of our proposed region-guided decoder with anchor pair (static anchor and dynamic anchor), where PE means positional encoding.
@@
where PE($\cdot$) means positional encoding to generate sinusoidal embeddings. For clarity, we use $A^{j}_{s}$ and $P^{j}_{s}$ to sign the static anchor and its positional embedding in $j$-th decoder layer, even though it is never updated. With the explicit initialization, regional priors are introduced to guide the decoder in generating non-overlapped proposals.
Anchor Pair Update.  Although introducing regional guidance by explicit initialization, maintaining the guidance during decoding iterations is also important. Following this idea, static anchors are designed to maintain guidance without updating, while dynamic anchors are designed to update for localization as shown in Fig. 3. For static anchors,
@@
Given dynamic anchors $A^{j}_{d}=(a_{c}^{j},a_{\sigma}^{j})$ in $j$-th decoder layer and the relative positions $\Delta A^{j}_{d}$ = ($\Delta a^{j}_{c},\Delta a^{j}_{\sigma}$) from a prediction head, the dynamic anchors are updated as:
$\begin{split}A^{j+1}_{d}=A^{j}_{d}+\Delta&A^{j}_{d}=(a_{c}^{j}+\Delta a^{j}_{c},\;a_{\sigma}^{j}+\Delta a^{j}_{\sigma}),\\ P^{j+1}_{d}&=\text{MLP}(\text{PE}(A^{j+1}_{d})).\end{split}$ (5)
Note that all prediction heads share the same parameters.